### v2

# Infinite-Horizon Optimal Wireless Control Over Shared State-Dependent Fading Channels for IIoT Systems

Abstract
Heterogeneous systems consisting of a multiloop wireless control system (WCS) and a mobile agent system (MAS) are ubiquitous in Industrial Internet of Things systems. Within these systems, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance, requiring joint coordination between the WCS and MAS. Such coordination introduces different time steps and hybrid state spaces consisting of logical components and continuous components. This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to safety constraints. A state-dependent fading channel is modeled to capture interference among transmission links, as well as the effects of mobile agents’ movements on successful wireless transmission. In order to address the heterogeneous system dynamics, the optimal control problem is formulated as the optimal constrained set stabilization of the MAS by establishing a necessary and sufficient condition for the Lyapunov-like performance of WCS with the expected decay rates. Using the semi-tensor product of matrices, a constrained optimal state transition graph is constructed to encode the constrained system dynamics as well as objective function, which further reduces the problem into a minimum-mean cycle problem for the graph. By studying the properties of the graph, the feasibility is proven, and an effective algorithm is proposed for the construction of optimal input sequences. An illustrative example is provided to demonstrate effectiveness of the proposed method.
keywords:
Heterogeneous system; infinite-horizon optimal control; shared wireless medium; constrained set stabilization; semi-tensor product of matrices.
thanks: The work was supported by National Key R&D Program of China under the grant 2022YFB3303900 and the NSF of China under the grants 61933009, 62103272. This paper was not presented at any IFAC meeting. Corresponding author S. Zhu.
1 Introduction
1.1 Background and motivation
Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields, including smart manufacturing, through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011]. These plants consist of multiple physical systems with sensing, computation, communication and actuation capabilities, which flexibly interact with one another and human operators [Ahlen2019, vitturi2019]. Due to such complex interactions among the entities, wired communication will reach its limits. Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming multiloop wireless control systems (WCSs).
An integral aspect of IIoT in smart manufacturing is the use of mobile agents, e.g., automated guided vehicles (AGVs) [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019]. In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as moving within task areas and velocity limits while avoiding dangerous positions [Baumann2021]. Moreover, the positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over [Agrawal2014, Quevedo2013]. On-site measurements in factories reveal that the channel fading caused by machinery movements exhibits a piecewise behavior, i.e., the fading channel gain remains stationary as the machinery moves within a certain area [Ahlen2019]. And each area is associated with a certain level of shadowing effects. This corresponds to a high level dynamics of MAS exhibiting logical characteristics. Additionally, the movements of mobile agents among areas after spatial discretization is typically on the time scale of seconds, while the wireless channel transmits packets on the order of milliseconds. This means that the coordination of the WCS and mobile agent system (MAS) introduces different time steps and hybrid state spaces, consisting of the continuous component of industrial plants and the logical component of mobile agents. These distinctive features make the modeling and analysis tools on traditional WCSs fail in addressing the IIoT system considered in the paper. Due to the complex interactions between the WCS and MAS, in order to ensure the control performance of the WCS, it is necessary to explicitly examine the influence of mobile agents’ positions on the channel states of the WCS, and to design a constrained motion strategy for agents such that the desired performance requirements are guaranteed for the WCS.
1.2 Related work
Since the typical connection density in the IIoT scenario is $10^{6}/km^{2}$, transmission scheduling among sensors over the shared limited wireless medium is an important topic in practice [valerio2021]. For example, static schedules typically specifying periodic communication sequences were proposed to satisfy given control performance requirements such as stability [Hristu2001], controllability and observability [zhang2006] and minimizing linear quadratic objectives [le2011]. Dynamic schedules deciding access to the shared wireless medium dynamically at each step were also extensively studied [donkers2011, gatsis2015, Molin]. For a survey on design and optimization for WCSs, please refer to [Park2018].
As important components of IIoT systems, the MAS and WCS are typically considered separately in the literature. Recently, the finite-field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014]. The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011]. It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints. Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems. Noting that all the above results are established for finite-valued systems, i.e., the dynamics of all nodes in the system are the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to control performance requirements constraints. As for the other part of IIoT, WCSs with fading channels have been extensively studied in the past few decades. It is noted that the channel gains used to characterize shadow fading are traditionally modeled as independent identical distributed random processes [Gatsis2014] or Markov chains [zhangq1999], where the network state is assumed to be independent from physical states. With such independency, several results were established for WCSs over fading channels including co-design problem for communication and control systems [Gatsis2014], state estimation [Quevedo2012] and so on. References [gatsis2015] and [gatsis2018] considered shared fading channels, where centralized and distributed channel-aware access mechanisms were respectively presented to meet the Lyapunov-like performance requirements while minimizing the total expected power consumption. The above channel models are clearly inadequate for characterizing the complex interactions between the WCS and MAS. In addition, all these results focus on the control performance of plants with only continuous state space. It is obvious that these methods fail in analyzing IIoT systems with hybrid state spaces considered in this paper.
Few works addressing dependence of network state on physical states have been conducted in IIoT. A state-dependent bursty fading channel model was proposed in [hu2014]. However, only homogeneous systems without constraints were considered in these works. In order to model random time variations of wireless environment, [Quevedo2013] introduced a network state process for state estimation, but the network topology was assumed to be fixed. The optimal network topology configuration was then determined for each network state to minimize an expected error covariance measure [Leong2016]. However, the external environment in [Leong2016, Quevedo2013] was modeled as a (semi)-Markov chain which cannot be controlled. Removing the uncontrollability assumptions and modeling the external environment as a Markov decision process, [hu2019] studied the infinite-horizon optimal design problem of communication and control, such that the single-loop WCS with state-dependent channel achieves both safety and efficiency. By formulating this problem as a constrained two-player cooperative game, approximate optimal control and transmission power policies were obtained by solving relaxed convex generalized geometric programs. Similar problem was studied in [hu2022], which considered a generalized state-dependent Markov channel model to address time varying data rates. By solving a constrained polynomial optimization problem, optimal design strategies were obtained. In the above works, a common time step is adopted for the evolution of the MAS and WCS. Noting that the wireless channel typically transmit packets on the order of milliseconds, this requires mobile agents to move on a millisecond time scale, which is often unrealistic. In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, the state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
1.3 Contributions
In this paper, we focus on the optimal control of MAS with safety constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling. A preliminary version of the paper appears in [wangs2022], where a common time step is considered for the WCS and the leader-follower MAS. The main contributions of the paper are summarized as follows.
  • We propose a heterogeneous IIoT system framework to capture the coupling between the WCS and MAS, where the multiloop WCS is described as a switched dynamics model depending on wireless transmission, and the finite-field architecture is used to model the high level dynamics of MAS moving among areas with different levels of shadow effects. Different time steps are adopted to capture different time scales between the wireless transmission of control systems and the movements of mobile agents among discretized areas. Effect of mobile agents’ movements on successful wireless transmission is characterized by a state-dependent fading channel, which captures the spectrum resource competition and interference among different loops. Based on this model, the infinite-horizon optimal control problem is established for the heterogeneous IIoT system to quantify the tradeoff between WCS’s power consumption and MAS’s control cost, while ensuring the Lyapunov-like performance requirement of the WCS and the safety constraints of the MAS.
  • To address the hybrid nature of the heterogeneous IIoT system, we first equivalently convert the optimal control problem into the optimal constrained set stabilization problem of the MAS. In order to address the nonlinearity due to the finite-field model of the MAS, we adopt the semi-tensor product approach, and construct a constrained optimal state transition graph to handle safety constraints elegantly. By encoding the constrained system dynamics and objective function into the graph, the problem is further deduced into a minimum-mean cycle problem on the constrained optimal state transition graph. Then, through studying the properties of the graph, we propose an efficient algorithm to design optimal input sequences that ensure wireless control performance and safety constraints.
This paper is organized as follows. In Section 2, we characterize the heterogeneous IIoT system and formulate the problem studied in this paper. Section 3 formulates the optimal control problem as the optimal constrained set stabilization problem for the MAS. Design of optimal controllers using the graphical method is provided in Section 4. Section 5 demonstrates main results by an illustrative example, which is followed by the conclusion in Section 6. The key notations used in this paper are summarized in Table 1. Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$, and $\otimes$ is the Kronecker product of matrices [chengqi2011, cheng2011]. Since the semi-tensor product is a generalization of the conventional matrix product, the symbol $\ltimes$” is omitted in most places of this paper when no confusion arises.
Table 1: Notations.
Notations Definitions
$\mathcal{D}_{s}$ Logic domain $\{0,1,\cdots,s-1\}$
$I_{s}$ $s$-dimensional identity matrix
$\delta_{s}^{i}$ $i$-th column of $I_{s}$
$[F]_{:,j}$ $j$-th column of matrix $F$
$[F]_{i,j}$ $(i,j)$-th entry of matrix $F$
$\mathbb{R}^{n\times m}$ Set of $n\times m$ real matrices
$Col(F)$ Set $\{[F]_{:,j}:j=1,\cdots,m\}$ for $F\in\mathbb{R}^{n\times m}$
$\ltimes$ Semi-tensor product
[…]
${}\mathbb{P}\{\cdot\}$ Probability of a random event
$\mathbb{E}\{\cdot\}$ Expectation of a random variable
$|\mathcal{M}|$ Cardinal number of set $\mathcal{M}$
2 System model and problem formulation
We consider a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a typical smart manufacturing scenario (Fig. 1). In the wireless control architecture, $q$ independent plants are controlled over a shared wireless medium, where the measurements of plant $i$, $i\in\{1,\cdots,q\}$ are wirelessly transmitted to the access point by sensor $i$ to compute the control inputs. At the same time, $n$ mobile agents perform manufacturing tasks in the same workspace. In this process, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over, resulting in the coupling between the WCS and MAS.
Figure 1: Framework of a heterogeneous IIoT consisting of a multiloop WCS and an MAS in a factory floor. ‘S’ and ‘A’ respectively represent sensing and actuation capabilities. The workspace is partitioned into $\kappa$ cells $\{0,1,\cdots,\kappa-1\}$, and each cell is abstracted as a logical state indicating a level of shadowing effects. Each mobile agent $j$ moves within its task area, leading to logical state transitions among cells in the state constraint set $\mathcal{C}_{\alpha_{j}}$.
2.1 Plants
Denote the state of plant $i$ at time instant $t_{l}$ by $x_{i}(l)\in\mathbb{R}^{n_{i}}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
$x_{i}(l+1)=A_{i}x_{i}(l)+B_{i}v_{i}(l)+\xi_{i}(l),$ (1)
where $\{t_{l}\}_{l=0}^{\infty}$ with $t_{l}<t_{l+1}$ is the sequence of sampling time instants, $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}}$ is the driving input, and $\{\xi_{i}(l):l=0,1,\cdots\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
2.2 Sensors and access point
$q$ sensors are placed to monitor the plants. At every time instant $t_{l}$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not. We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at time instant $t_{l}$ is $\gamma_{i}(l)\in\mathcal{D}_{r}$. Assume that the wireless communication policy111 Regardless of the effect of mobile agents’ movements on wireless transmission, following the approach in [gatsis2018], it can be proven that the optimal transmission policies ensuring the Lyapunov-like control performance of the WCS only depend on the channel states $\gamma_{i}(l)$. is pre-designed as
$\varphi_{i}(l)=h_{i}(\gamma_{i}(l)),$ (2)
where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(l)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at time instant $t_{l}$, and each sensor transmits only when its corresponding local channel state is favorable enough and backs off otherwise. In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the local channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot. The pilot signal may also serve as a synchronization purpose [gatsis2018, hu2011].
When sensor $i$ decides to transmit, it wirelessly transmits measurements of plant $i$ to the access point with a fixed power $\mu_{i}>0$, where the access point plays the role of remote controller responsible for computing control decisions based on the received measurements [ma2018]. Then, control decisions are wirelessly fed back to actuator $i$. These $q$ control loops form a multiloop WCS.
2.3 Multiloop WCS
The transmission of measurements and control decisions might fail due to packet decoding errors and packet collisions. Assume that both wireless uplink and downlink have the same channel characteristics [hu2019]. In this case, we focus on the uplink transmission from the sensors to the access point, and the idea can be extended into the downlink transmission scenario straightforwardly. Use $\lambda_{i}(l)\in\mathcal{D}_{2}$ to indicate the successful or fail transmission of sensor information for plant $i$ at time instant $t_{l}$. Then, the measurement $x_{i}(l)$ is used to compute the control input $v_{i}(l)$ at a successful transmission ($\lambda_{i}(l)=1$), leading to the asymptotically stable closed-loop system matrix $A_{c,i}$. Otherwise, an open-loop system with the system matrix $A_{o,i}$, which may be unstable, is obtained. The resulting system can be described as the following switched model:
$x_{i}(l+1)=\left\{\begin{array}[]{ll}A_{c,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=1,\\ A_{o,i}x_{i}(l)+\xi_{i}(l),&\lambda_{i}(l)=0,\end{array}\right.~{}i=1,\cdots,q.$ (3)
2.4 Finite-field MAS
An MAS coordinates with the multiloop WCS to jointly perform smart manufacturing tasks. Mobile agents are driven by the standard single-integrator dynamics222 This paper uses the model as an illustration of how to abstract the finite-field MAS from the physics dynamics. The proposed results are general and can be extended to other models for mobile agents. as
$z_{j}(k+1)=z_{j}(k)+T\nu_{j}(k),j=1,\cdots,n,$
where $z_{j}(k),\nu_{j}(k)\in\mathbb{R}^{2}$ respectively denote the position and velocity of agent $j$ in the two-dimensional space at time instant $t_{\tau k}$, $\{t_{\tau k}\}_{k=0}^{\infty}$ is the sequence of sampling time instants with time step $T$, and $\tau$ is a positive integer indicating different time steps for the evolution of the MAS and WCS.
When coordinating with the WCS to jointly perform overall tasks, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over. Note that what matters for shadow fading is the position in space within certain intervals [Ahlen2019], not the specific position determined by the single-integrator dynamics. Partition the whole workspace into $\kappa$ non-overlapping cells $\{0,1,\cdots,\kappa-1\}$. Abstract the cell in which each mobile agent is located as a logical state in $\mathcal{D}_{\kappa}$. Each agent interacts with others by on-board sensors, and updates its state as a weighted combination of its previous state, those of its in-neighbors and the input, determined by the following finite-field model:
$\alpha_{j}(k+1)\!=\!\{_{\kappa}\!\sum\limits_{l\in\mathcal{I}_{j}\cup\{j\}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)+_{\kappa}u_{j}(k),j=1,\cdots,n,$ (4)
where $\mathcal{I}_{j}$ is the set of all in-neighbors of $j$, $a_{j,l}\in\mathcal{D}_{\kappa}$ denotes the weight, operations $+_{\kappa}$ and ×$\times_{\kappa}$ respectively denote the modular addition and modular multiplication over $\mathcal{D}_{\kappa}$ [Lidl1996], $\alpha_{j}(k)$, $u_{j}(k)\in\mathcal{D}_{\kappa}$ respectively denote the state and control input of agent $j$ at time instant $t_{\tau k}$. The feedback term $\{_{\kappa}\sum_{l\in\mathcal{I}_{j}}\}a_{j,l}\times_{\kappa}\alpha_{l}(k)$ together with $u_{j}(k)$ can describe several mobile agents’ controllers, including consensus, rendezvous and formation [Mesbahi, Pasqualetti2014, Sundaram2013]. The constraint on state $\alpha(k)=(\alpha_{1}(k),\cdots,\alpha_{n}(k))\in\mathcal{C}_{\alpha}\subseteq\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}$ with $\alpha_{j}(k)\in\mathcal{C}_{\alpha_{j}}$, $j=1,\cdots,n$ is imposed to ensure each agent $j$ moves within its task areas $\mathcal{C}_{\alpha_{j}}$ while avoiding dangerous positions, such as collisions with industrial plants. In addition, the transition constraint $u(k)=(u_{1}(k),\cdots,u_{n}(k))\in\mathcal{C}_{u}(\alpha(k))$ is imposed to restrict agents’ movements within the actual velocity limits. Such safety constraints distinguish the model from previous works [hu2022, hu2019] representing the mobile agent by an unconstrained Markov decision process.
2.5 State-dependent fading channel model
In the considered IIoT system, movements of agents may lead to shadow fading on the wireless channel between sensors and the access point, and finally influence the probability of successful wireless transmission. With different time steps, during each sampling interval $[t_{\tau k},t_{\tau(k+1)})$ of the MAS, there are $\tau$ time instants of the wireless transmission as $t_{l}$, $l={\tau k},{\tau k+1},\cdots,{\tau(k+1)-1}$, where the $k$ is determined by the relation $k=\lfloor l/\tau\rfloor$. This means that the successful transmission probabilities for WCS during these $\tau$ consecutive time instants $t_{l}$ depend on the state of MAS at time instant $t_{\tau k}$, i.e., $\alpha(k)$. Then, the WCS and MAS coupling is modeled by the following state-dependent fading channel:
$\mathbb{P}\Big{\{}\lambda_{i}(l)=1|\alpha(k)=\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}=\bar{\lambda}_{i}(\alpha),$ (5)
where $\lambda_{i}(l)=1$ indicates the successful transmission of link $i$ at time instant $t_{l}$.
Let $\eta_{i}(l)=1$ denote the successful packet decoding of link $i$. As shown in Fig. 2, with wireless communication policy (2), it holds
$\bar{\lambda}_{i}(\alpha)\!=\!\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)\mathbb{P}\{\eta_{i}(l)\!=\!1|\alpha(k)\!=\!\alpha,k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\},$
where $\bar{\gamma}_{i}(c,\alpha)=\mathbb{P}\{\gamma_{i}(l)=c|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$ characterizes the probability distribution for the local channel state, and can be easily obtained by measuring the local fading amplitude for each MAS state by the statistical methods in [Agrawal2014, Kashiwagi2010]. Taking into account the interference among transmission links by adopting the collision channel model [gatsis2018], the last term in $\bar{\lambda}_{i}(\alpha)$ is equal to $\mathbb{P}\{\hbox{SNR}_{i}(l)\geq\eta_{i}^{0}|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor,\varphi_{i}(l)=1,\varphi_{j}(j)=0,j\neq i\}\prod_{j\neq i}\mathbb{P}\{\varphi_{j}(l)=0|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}$, where $\hbox{SNR}_{i}(l)$ and $\eta_{i}^{0}$ denote the signal-to-noise ratio and the specified decodability threshold, respectively.
Figure 2: State-dependent fading channel (5).
Eq. (5) presents a general state-dependent channel model. In this model, the probability of a successful transmission for WCS (3) depends on the state of MAS (4). In addition, as shown in Fig. 2, Eq. (5) takes into account wireless communication policies, the interference among transmission links, and different time steps for the WCS and MAS, which cannot be captured by the existing channel models in [hu2022, hu2019].
2.6 Problem formulation
Considering the state-dependent fading channel model (5), given an initial state $\alpha_{0}\in\mathcal{C}_{\alpha}$, our primary goal is to design an input sequence ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ for MAS (4) such that $\alpha(k)\in\mathcal{C}_{\alpha}$, $k=0,1,\cdots$, and under ${\bf{u}}$ the Lyapunov-like performance of the WCS is satisfied, which is formulated as follows.
Definition 1.
The Lyapunov-like performance is said to be guaranteed for WCS (3) with the state-dependent fading channel (5), if there exist an input sequence ${\bf{u}}$ and a positive integer $L_{0}$ such that the inequalities
$\mathbb{E}\Big{[}V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\leq\rho_{i}V_{i}(x_{i}(l))+Tr(Q_{i}\Xi_{i}),\forall x_{i}(l)\in\mathbb{R}^{n_{i}}$ (6)
hold for any $l\geq L_{0}$, any $x_{i}(l)\in\mathbb{R}^{n_{i}}$ and any $i\in\{1,\cdots,q\}$, where $V_{i}(x_{i})=x_{i}^{\top}Q_{i}x_{i}$ is a given quadratic Lyapunov function with $Q_{i}$ being an $n_{i}\times n_{i}$ real positive definite matrix and $\rho_{i}\in(0,1)$, $i=1,\cdots,q$.
In (6), $\rho_{i}$ represents the desired decay rate and the trace $Tr(Q_{i}\Xi_{i})$ denotes a persistent noise perturbation. Similar kinds of performance requirements as in (6) have been adopted for WCSs [gatsis2015, gatsis2018]. Different from them, Definition 1 explicitly explores the effect of mobile agents’ movements on the successful wireless transmission of the WCS. In addition, it allows adopting different time steps to capture different time scales between the wireless transmission of control systems and the movements of mobile agents from cell to cell.
Our second goal in this paper is to determine optimal input sequence ${\bf{u}}^{\ast}$ such that the average cost function
$J_{{\bf{u}}}=\lim\limits_{K\rightarrow\infty}\frac{1}{K}\Big{\{}\sum\limits_{l=0}^{K-1}\sum\limits_{i=1}^{q}\mu_{i}\mathbb{P}\Big{\{}\varphi_{i}(l)=1|\alpha(k),k=\Big{\lfloor}\frac{l}{\tau}\Big{\rfloor}\Big{\}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\lambda\sum\limits_{k=0}^{\lfloor(K-1)/\tau\rfloor}g(\alpha(k),u(k))\Big{\}}$ (7)
reaches optimality for the initial state $\alpha_{0}$, where $\lambda$ is a positive constant, $\mu_{i}$ is the transmit power for sensor $i$, $\sum_{i=1}^{q}\mu_{i}\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$ quantifies the power consumption of the WCS, $g:\mathcal{D}_{\kappa}\times\cdots\times\mathcal{D}_{\kappa}\rightarrow\mathbb{R}_{+}$ is the stage cost function of the MAS, and $\mathbb{R}_{+}$ is the set of positive real numbers, i.e., to solve the infinite-horizon optimal control problem
$P_{1}:$ $\min_{{\bf{u}}}J_{{\bf{u}}}$
[…]
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(0)=\alpha_{0},$
$~{}~{}~{}~{}~{}~{}~{}~{}\alpha(k)\in\mathcal{C}_{\alpha},u(k)\in\mathcal{C}_{u}(\alpha(k)),k=0,1,\cdots.$ (8)
3 Optimal wireless control as optimal constrained set stabilization
In the optimal control problem $P_{1}$, there are heterogeneous system dynamics constraints consisting of continuous component (3) and logical component (4) with different time steps. In addition, control performance (6) of the continuous system is constrained by the logical one. The disparity in analyzing these distinct systems makes the resolution of the problem difficult. To address these challenges, we convert WCS’s performance requirements (6) together with safety constraints (8) into the constrained set stabilization of MAS (4) using the state-dependent fading channel (5). In this way, problem $P_{1}$ is converted to the optimal constrained set stabilization problem of the MAS.
For the convenience of statement, we call ${\bf{u}}=\{u(k)\in\mathcal{C}_{u}(\alpha(k)):k=0,1,\cdots\}$ an admissible input sequence in the following.
Definition 2.
Given $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$, MAS (4) is constrained $\mathcal{M}$-stabilizable from initial state $\alpha_{0}$, if there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$ holds for any $k<K_{0}$, while $\alpha(k)\in\mathcal{M}$ holds for any $k\geq K_{0}$.
Definition 3.
A subset $\mathcal{M}\subseteq\mathcal{C}_{\alpha}$ is called a constrained control invariant subset of MAS (4), if for any $\alpha_{0}\in\mathcal{M}$, there exists an admissible input sequence ${\bf{u}}$ such that $\alpha(k)\in\mathcal{M}$ holds for any $k=0,1,\cdots$.
Obviously, the union of any two constrained control invariant subsets of MAS (4) is another constrained control invariant subset. The union of all constrained control invariant subsets contained in $\mathcal{M}$ is called its largest constrained control invariant set, denoted by $I(\mathcal{M})$.
Theorem 4.
The Lyapunov-like performance (6) is guaranteed for WCS (3) with the state-dependent fading channel (5) and safety constraints (8), if and only if MAS (4) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, where $\Omega(s)=\{\alpha\in\mathcal{C}_{\alpha}:\bar{\lambda}_{i}(\alpha)\geq s_{i},i=1,\cdots,q\}$, $s=[s_{1}~{}\cdots~{}s_{q}]$ with
$s_{i}=\sup\limits_{y\in\mathbb{R}^{n_{i}},y\neq 0}\frac{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})y}{y^{\top}(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})y}.$ (9)
Proof 3.1.
For each subsystem $i$, since $\mathbb{E}[\xi_{i}(l)]=\mathbf{0}$ and $\lambda_{i}(l)$ is independent of $x_{i}(l)$, by (3), it holds
$\mathbb{E}[V_{i}(x_{i}(l+1))|x_{i}(l),\alpha(k),k=\lfloor l/\tau\rfloor]$
$=$ $\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{c,i}^{\top}Q_{i}A_{c,i}x_{i}(l)$
+$+\mathbb{P}\{\lambda_{i}(l)=0|\alpha(k),k=\lfloor l/\tau\rfloor\}x_{i}^{\top}(l)A_{o,i}^{\top}Q_{i}A_{o,i}x_{i}(l)$
$+T_{r}(Q_{i}\Xi_{i}).$
Then, condition (6) subject to multiloop WCS dynamics (3) and safety constraints holds, if and only if there exist an admissible input sequence and a positive integer $L_{0}$ such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $x_{i}(l)\neq\mathbf{0}$, it holds
$\mathbb{P}\{\lambda_{i}(l)=1|\alpha(k),k=\lfloor l/\tau\rfloor\}$
$\geq$ $\frac{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-\rho_{i}Q_{i})x_{i}(l)}{x_{i}^{\top}(l)(A_{o,i}^{\top}Q_{i}A_{o,i}-A_{c,i}^{\top}Q_{i}A_{c,i})x_{i}(l)},$ (10)
which are equivalent to the condition that there exist an admissible input sequence ${\bf{u}}$ and a positive integer $K_{0}$, such that $\alpha(k)\in\mathcal{C}_{\alpha}$, and for all $k\geq K_{0}$, it holds $\bar{\lambda}_{i}(\alpha(k))\geq s_{i}$, $s_{i}$ is given in (9), that is, MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$.
Next, we prove that MAS (4) is constrained $\Omega(s)$-stabilizable from initial state $\alpha_{0}$ if and only if it is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$. We just need to prove the necessity. This is proved by a contradiction argument. Since MAS (4) is constrained $\Omega(s)$-stabilizable from $\alpha_{0}$, we can obtain a feasible admissible input sequence ${\bf{u}}^{\prime}$ and a positive integer $K_{0}$, such that for any integer $k<K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\mathcal{C}_{\alpha}$; for any integer $k\geq K_{0}$, it holds $\alpha(k;\alpha_{0},{\bf{u}}^{\prime})\in\Omega(s)$.
Assuming that MAS (4) is not constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then under input sequence ${\bf{u}}^{\prime}$, there exists an integer $k_{0}\geq K_{0}$ satisfying $\alpha(k_{0})\notin I(\Omega(s))$. Since $I(\Omega(s))$ is a largest constrained control invariant subset, for any $\alpha^{\prime}\in\Omega(s)$ and any admissible input sequence ${\bf{u}}$, if $\alpha^{\prime}\notin I(\Omega(s))$, there exists an integer $k(\alpha^{\prime},{\bf{u}})$ such that $\alpha(k(\alpha^{\prime},{\bf{u}}))\notin\Omega(s)$ from $\alpha^{\prime}$. Then, under input sequence ${\bf{u}}^{\prime}$, it holds $\alpha(k_{0}+k(\alpha(k_{0}),{\bf{u}}^{\prime\prime}))\notin\Omega(s)$ from $\alpha_{0}$, $u^{\prime\prime}(k)=u^{\prime}(k+k_{0})$, which is a contradiction. Thus, MAS (4) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$. $\Box$
By (10), $s_{i}$ is the lower bound of the successful transmission probability for system $i$ ensuring the desired Lyapunov-like performance in Definition 1. The value of $s_{i}$ can be obtained by the following linear program:
$\min_{\theta_{i}\geq 0}~{}\theta_{i}$
[…]
which can be solved using the toolbox such as YALMIP in MATLAB. In addition, $\Omega(s)$ consists of all constrained states of the MAS satisfying the performance requirements constraint (6) of WCS (3).
For the optimal control problem $P_{1}$, all constraints can be guaranteed if MAS (4) is constrained $I(\Omega(s))$-stabilizable. Since $\mathbb{P}\{\varphi_{i}(l)=1|\alpha(k)=\alpha,k=\lfloor l/\tau\rfloor\}=\sum_{c=0}^{r-1}\bar{\gamma}_{i}(c,\alpha)h_{i}(c)$, ${\bf{u}}^{\ast}$ is a solution to $P_{1}$, if and only if it is a solution to the following infinite-horizon optimal constrained set stabilization problem of MAS (4):
$P_{2}:$ $\min_{{\bf{u}}}\bar{J}_{{\bf{u}}}=\lim_{K\rightarrow\infty}\frac{1}{K}\sum_{k=0}^{K-1}\bar{g}(\alpha(k),u(k))$
[…]
$~{}~{}~{}~{}~{}~{}~{}\,\exists\ K_{0}>0,\alpha(k)\in I(\Omega(s)),\forall k\geq K_{0},$
$~{}~{}~{}~{}~{}~{}~{}\,\alpha(k)\in\mathcal{C}_{\alpha},\forall k<K_{0},$
where $\bar{g}(\alpha(k),u(k))$ is the joint stage cost defined as
$\bar{g}(\alpha(k),\!u(k))\!=\!\tau\!\sum_{i=1}^{q}\!\mu_{i}\!\sum_{c=0}^{r-1}\!\bar{\gamma}_{i}(c,\alpha(k))h_{i}(c)\!+\!\lambda g(\alpha(k),\!u(k)).$
Let the optimal objective values of problems $P_{1}$ and $P_{2}$ be $J^{\ast}=J_{{\bf{u}}^{\ast}}$ and $\bar{J}^{\ast}=\bar{J}_{{\bf{u}}^{\ast}}$, respectively. It is easy to obtain that $J^{\ast}=\bar{J}^{\ast}/\tau$.
4 Infinite-horizon optimal controller design
Note that both the state space and control space of the MAS dynamics (4) are finite. This property allows us to adopt the graphical method to address the optimization problem $P_{2}$. By reformulating the logical-state MAS in a tractable form via the semi-tensor product approach, the constrained optimal state transition graph is firstly constructed to encode the constraints and the objective function of $P_{2}$. Based on this, the feasibility of problem $P_{2}$ is deduced to the existence of cycles in the graph, and the design of optimal input sequences is modeled as a minimum-mean cycle problem. An algorithm is then proposed to construct optimal controllers.
4.1 Feasibility analysis via a graphical method
Due to the highly nonlinear finite-field model, it is difficult to apply classical nonlinear system theory to investigate (4). Recently, the semi-tensor product approach has emerged as an elegant tool to deal with such issues [chengqi2011, cheng2011]. The basic idea is to embed the nonlinear finite-filed dynamics in (4) into a higher dimensional space where its evolution is linear. Specially, representing $i-1\in\mathcal{D}_{\kappa}$ as the vector $\delta_{\kappa}^{i}$, $i=1,\cdots,\kappa$, then $\alpha_{i}(k),u_{i}(k)\in\mathcal{D}_{\kappa}$ have their corresponding vector forms (still using the same symbols) $\alpha_{i}(k),u_{i}(k)\in Col(I_{\kappa})$, respectively. According to similar calculation to [wangs2022], we convert MAS (4) into the following bilinear equivalent algebraic form:
$\alpha(k+1)=Fu(k)\alpha(k),$ (11)
where $\alpha(k)=\ltimes_{i=1}^{n}\alpha_{i}(k)\subseteq Col(I_{N})$, $u(k)=\ltimes_{i=1}^{n}u_{i}(k)\subseteq Col(I_{N})$, $N:=\kappa^{n}$, and $F$ is a logical matrix [chengqi2011, cheng2011] characterizing all state-transition information of the MAS. Then, the set of admissible inputs steering MAS (11) from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ in one step can be obtained as
$U_{a,b}=\{\delta_{N}^{l}\in\mathcal{C}_{u}(\delta_{N}^{a}):[F\delta_{N}^{l}]_{b,a}=1\}.$ (12)
In addition, the set of optimal inputs enabling the transition with the lowest cost is
$\bar{U}_{a,b}={\arg\min}_{\delta_{N}^{l}\in U_{a,b}}\bar{g}(\delta_{N}^{a},\delta_{N}^{l}).$ (13)
Figure 3: Constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$.
To determine the feasibility of optimization problem $P_{2}$, we construct the constrained optimal state transition graph $G=(\mathcal{V},\mathcal{E},w)$ to encode constraints and the objective function. Regard each state $\delta_{N}^{a}\in\mathcal{C}_{\alpha}$ as a vertex on a directed graph, and encode each constrained state transition from $\delta_{N}^{a}$ to $\delta_{N}^{b}$ into edge $(\delta_{N}^{a},\delta_{N}^{b})$. Additionally, assign the weight $w(\delta_{N}^{a},\delta_{N}^{b})=\bar{g}(\delta_{N}^{a},\delta_{N}^{l})$ to the edge $(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}$ to encode the lowest joint cost under the optimal input $\delta_{N}^{l}\in\bar{U}_{a,b}$. See Fig. 3 for an illustration.
For an input sequence ${\bf{u}}$, we say that the state trajectory steered by it converges to a cycle $\mathcal{C}$, if the path $\mathcal{T}=\{\alpha(k):k=0,1,\cdots\}$ in $G$ keeps repeating $\mathcal{C}$ after a finite time. Define the weight and average weight of a path $\mathcal{T}=\{\delta_{N}^{a_{0}},\cdots,\delta_{N}^{a_{l}}\}\subseteq\mathcal{V}$, $l\geq 0$ as $w(\mathcal{T})=\sum_{i=0}^{l-1}w(\delta_{N}^{a_{i}},\delta_{N}^{a_{i+1}})$ and $\bar{w}(\mathcal{T})=w(\mathcal{T})/l$, respectively. Then, we have the following result.
Theorem 5.
MAS (11) is constrained $I(\Omega(s))$-stabilizable from initial state $\alpha_{0}$, if and only if $I(\Omega(s))\cap\mathcal{R}(\alpha_{0})\neq\emptyset$, where $\mathcal{R}(\alpha_{0})$ denotes the set of constrained states reachable from $\alpha_{0}$. In addition, if MAS (11) is constrained $I(\Omega(s))$-stabilizable from $\alpha_{0}$, then
  1. (i)
    it holds $C(G[\Phi])\neq\emptyset$, where $\Phi=I(\Omega(s))\cap\mathcal{R}(\alpha_{0})$, $G[\Phi]=(\Phi,\mathcal{E}^{\prime},w)$ is the induced subgraph of $G$ induced by $\Phi$ with =$\mathcal{E}^{\prime}=\{(\delta_{N}^{a},\delta_{N}^{b})\in\mathcal{E}:\delta_{N}^{a},\delta_{N}^{b}\in\Phi\}$, and $C(\cdot)$ represents the set of all cycles in the graph;
  2. (ii)
    solutions to the infinite-horizon optimal control problem $P_{2}$ exist.